# Investigating Length Issues in Document-level Machine Translation

### v1

@@ 5.2 Models
We used the unifPE algorithm to fine-tune two pre-trained MT systems that were not trained with TED talks. As unifPE is designed for APEs, we considered NLLB200-distilled-600M [^13] https://huggingface.co/facebook/nllb-200-distilled-600M or NLLB for short (Costa-jussà et al., 2024) as a representative encoder-decoder model based on APEs. NLLB is a 12-layer encoder-decoder multilingual MT model pre-trained in 200 languages. We used the HuggingFace implementation which relies on sinusoidal APEs (Vaswani et al., 2017). We denote MT systems trained on TED-U (resp. TED-G) as FT-NLLB-U (resp. FT-NLLB-G), and Unif-NLLB-U (resp. Unif-NLLB-G) when fine-tuning with unifPE.
We also experiment with an LLM-based architecture, TowerBase-7B [^14] https://huggingface.co/Unbabel/TowerBase-7B-v0.1 (Alves et al., 2024) (TowerBase for short), derived from Llama2 using translation-related tasks. TowerBase uses RoPE (Su et al., 2024) to encode RPEs. As mentioned by Peng et al. (2024), they nonetheless encode some form of APE signal in some dimensions, and may therefore be also mildly impacted by the PE training distribution. We denote FT-Tower-U (resp. FT-Tower-G) and Unif-Tower-U (resp. Unif-Tower-G) the model fine-tuned on TED-U (resp. TED-G) without or with unifPE.
Both models were pretrained with large amounts of EN–FR data; we focus exclusively on the EN into FR direction. disabledisabletodo: disableCheck directionDetails on fine-tuning and decoding parameters can be found in Appendix A.4.
@@ 6.1 Length Issues
We report the ds-BLEU (Table 2) and COMET (Appendix, Table 6) scores of the pretrained models NLLB and TowerBase for multiple testsets, varying the average input segment lengths from one sentence to the maximum input length used in training. [^15] As explained in Section 3.2, these COMET scores require the realignment of target sentences with the reference.disabledisabletodo: disableCOMET scores in appendix ? For NLLB, we observe a drop of around $10$ ds-BLEU points and about $0.2$ COMET scores when translating test sets of $l_{max}=256$ instead of isolated sentences. Scores and their associated brevity penalties (BPs) only get worse with larger context lengths. For TowerBase, the decrease in BLEU is more progressive, with a sharp decline for all testsets for $l_{max}>1024$. The related COMET scores plummet immediately with the context size of $256$. Even though TowerBase is based on Llama2, which accepts inputs up to $4096$ tokens, the continued-pretraining mostly uses isolated sentences, which introduces an inductive bias affecting its ability to translate long texts.
@@
TED-U TED-G FT Unif
FT vs Unif FT vs Unif U vs G U vs G
sent 3.3 (0.00) 1.2 (0.00) - -2.1 (0.00)
256 - - - -0.6 (0.01)
512 -0.5 (0.01) - -0.7 (0.00) -0.4 (0.01)
768 -0.8 (0.00) -3.7 (0.00) 5.5 (0.00) 2.6 (0.00)
1024 - -7.8 (0.00) 12.2 (0.00) 5.1 (0.00)
1200 -2.3 (0.00) -8.5 (0.00) 12.7 (0.00) 6.5 (0.00)
1600 -2.6 (0.01) -8.8 (0.00) 11.8 (0.00) 5.6 (0.00)
2048 -3.3 (0.00) -7.3 (0.00) 10.7 (0.00) 6.8 (0.00)
sent 1.9 (0.00) 0.9 (0.00) - -0.9 (0.00)
256 - 0.4 (0.00) -0.8 (0.00) -0.5 (0.01)
512 -0.6 (0.02) - -0.4 (0.03) -
768 -0.7 (0.00) -4.7 (0.00) 11.2 (0.00) 7.2 (0.00)
1024 -2.2 (0.00) -7.5 (0.00) 19.3 (0.00) 14.0 (0.00)
1200 -5.3 (0.00) -6.5 (0.00) 15.7 (0.00) 14.5 (0.00)
1600 - -6.4 (0.00) 11.5 (0.00) 6.6 (0.00)
2048 - -4.7 (0.00) 10.4 (0.00) 7.2 (0.00)
Table 4: Average difference (and p-values) in ds-BLEU (top) evaluated on full TED talks and 100×100$\times$100 ×COMET (bottom) evaluated on realigned sentences for NLLB. Left: scores without and with unifPE; right: scores of fine-tuning on TED-U (U) or TED-G (G).
disabledisabletodo: disabletime permitting, make columns with even length
NLLB FT-U Unif-U FT-G Unif-G
$p_{0}$-$p_{1}$ 12.7 (0.00) - - 1.6 (0.00) 3.0 (0.00)
$p_{0}$-$p_{1}$ 7.2 (0.00) - - 1.9 (0.00) 2.4 (0.00)
$p_{2}$-$p_{3}$ 2.9 (0.00) 1.0 (0.00) -1.3 (0.00) 2.4 (0.00) -
$p_{3}$-$p_{4}$ 7.2 (0.00) 5.5 (0.00) 4.6 (0.00) 26.1 (0.00) 10.7 (0.00)
$p_{4}$-$p_{5}$ 3.9 (0.00) - 8.3 (0.00) 4.7 (0.00)
$p_{5}$-$p_{6}$ 3.3 (0.00) 31.6 (0.00) 27.1 (0.00) 6.1 (0.00) 15.4 (0.00)
Tower FT-U Unif-U FT-G Unif-G
$p_{0}$-$p_{1}$ 1.2 (0.02) - - - -
$p_{1}$-$p_{2}$ 0.6 (0.01) 0.6 (0.01)
$p_{2}$-$p_{3}$
$p_{3}$-$p_{4}$ 4.9 (0.00) 1.1 (0.00) 0.6 (0.04) 1.1 (0.00) 1.3 (0.00)
$p_{4}$-$p_{5}$ 1.7 (0.02) 1.4 (0.00) 2.1 (0.00) 2.0 (0.00) 1.8 (0.00)
$p_{5}$-$p_{6}$ 26.3 (0.00) 24.5 (0.00) 26.0 (0.00) 27.1 (0.00) 26.4 (0.00)
Table 5: Average difference (and p-values) of 100×100$\times$100 ×COMET-score evaluated on $794$ sentence pairs, translated at different positions by NLLB-based systems (top) and TowerBase-based systems (bottom).
@@
Again using the paired comparison methodology, we compare the performance of DLFT with and without unifPE. As shown in the left part of Table 4, fine-tuning using unifPE leads to steady improvements in translation scores for all test lengths, especially for systems fine-tuned with the unbalanced corpus (TED-G). The only exception is for sentence-level translations, which remain marginally better using standard DLFT than with unifPE. As Tables 11, 12, 13, 14 and 15 show, these improvements remain moderate, and the length issues continue to strongly impact translation scores, especially for test documents of $1024$ tokens or more.
For TowerBase, unifPE does not yield any significant difference with standard DLFT (see Appendix, Table 16), likely because this model relies on RPEs. From these comparisons, we conclude that unifPE partly resolves length issues for nllb, but hardly changes the situation for TowerBase.
@@
In the right part of Table 4, we evaluate the impact of the length distribution during fine-tuning for NLLB: the balanced distribution (TED-U) slightly but consistently underperforms the use of TED-G for short documents (fewer than $512$ tokens), a trend that is reversed for longer documents with strong improvement (over $768$ tokens). Manipulating the distribution of PEs with unifPE reduces the gap between the two fine-tuning corpora and makes the model more robust to document lengths rarely observed (or even unobserved) during fine-tuning. This analysis again reveals small differences between using ds-BLEU and COMET scores: in three cases out of $32$ comparisons (marked in bold), one metric detects a difference that is non-significant for the other.
@@ 6.2 Position Bias
To investigate potential translation issues related to large position indices, we collected the $794$ sentences that come from the final part of long talks and for which varying the window length also varied the position index. For each of them, we have seven translations, corresponding to positions $\{p_{0}^{j},\dots,p_{6}^{j}\},j\in\{1,\dots,794\}$. The average values for $\{p_{0}^{j},\dots,p_{6}^{j}\}$ are $\{p_{0},\dots,p_{6}\}=[66,173,262,335,585,779,1477]$. For this subset of sentences, we performed a paired t-test to compare the impact of the position on the translation score (using COMET as the only metric). We observe in Table 5 that in almost all comparisons but two, a small position index is preferable to a larger one. This suggests that one of the main challenges faced by Doc2Doc with large context lengths is to control the quality degradation for the final parts of the input text. Here again, unifPE slightly mitigates the problem for NLLB models, but no such improvement is observed for TowerBase.disabledisabletodo: disableWhy is tower different?
@@ 6.3 Repeated $n$-grams in Translation
One obvious problem with holistic translations produced by Nllb is the generation of outputs that are too short. A closer look at translation outputs also reveals that outputs contain many instances of repeated texts, usually occurring in the final part of the translation. To quantify this problem, we compute the percentage of translations of pseudo-documents in which the repetition of a long $n$-gram (with $n\geq 10$) is detected. Detailed results are in Appendix, Table 9: for all systems and fine-tuning strategies, the percentage of repetitions increases with the length, a problem that seems (for large text lengths) slightly more severe for TowerBase, which has a much better BP, than Nllb.
@@ 7 Conclusion
In this work, we have studied the ability of current MT architectures to handle long input texts, ideally entire documents, and to translate them holistically. Our analyses are based on systematic comparisons of translation outputs computed with varying input lengths, which are then evaluated with two automatic metrics. They consistently show that, even when the test document lengths match that of the training set and remain within the model limits, the translation scores tend to decrease with the source length, a degradation that mostly impacts sentences occurring far from the document start. We also show that manipulating the training distribution of lengths or PEs has a positive effect for APE-based models, which vanishes in RoPE-based models like TowerBase.disabledisabletodo: disableMake sure RPE is defined These results finally confirm the robustness of sentence-level baselines. They hint at the need to improve existing models to truly benefit from the potential of document-level MT, for instance by constraining the attention mechanism to simulate a form of sentence alignment, by improving the memorization capacities of existing architectures, or by ensuring that the generation algorithm does not eventually get trapped in repetition loops. These are some of the directions we wish to explore in future work.disabledisabletodo: disablePerhaps: more citations for this disable,color=cyan!30disable,color=cyan!30todo: disable,color=cyan!30Future: the errors beyond repetitions and missing material disabledisabletodo: disablethere should not be TEDtalks in NLLB training set, as the NLLB papers used IWSLT TED talks as benchmark to show the model performance, where is the only mention os TED talks.
@@ 8 Limitations
The empirical observations reported in this paper are only based on one single language direction,disabledisabletodo: disablecheck that this is clear and one domain (TEDtalks). This experimental design is motivated by (a) the fact that French-English is considered an easy pair for MT, with large sets of parallel training data; (b) TEDtalks data are a standard benchmark for document-level MT, and crucially contain very long parallel documents, allowing us to implement our evaluation methodology on a large range of length values. Furthermore, these datasets are not included in the training data of our models. We contend that the length issues observed in these favorable conditions for two representative systems would only be worse for more difficult or less-resourced language pairs.
@@ 10 Carbon Impact Statement
The experiments were conducted on a private infrastructure using a single A100 SXM4 GPU, with a carbon efficiency of $0.432$ kgCO2eq/kWh. The average time required for fine-tuning and checkpoint selection was $14.21$ hours for the four NLLB models, and $5.6$ hours for the Towerbase models. The average emissions are estimated to be $2.45$ kgCO2eq for NLLB-based models and $0.97$ kgCO2eq for models derived from TowerBase, with no offset applied. These estimations were based on the Machine Learning Impact calculator [^16] https://mlco2.github.io/impact#compute (Lacoste et al., 2019).

### v2

@@ 5.2 Models
We used the unifPE algorithm to fine-tune two pre-trained MT systems that were not trained with TED talks. As unifPE is designed for APEs, we considered NLLB200-distilled-600M [^15] https://huggingface.co/facebook/nllb-200-distilled-600M or NLLB for short (Costa-jussà et al., 2024) as a representative encoder-decoder model based on APEs. NLLB is a 12-layer encoder-decoder multilingual MT model pre-trained on 200 languages. We used the HuggingFace implementation, which relies on sinusoidal APEs (Vaswani et al., 2017). We also perform fine-tuning with SHAPE for comparison. We refer to the specific MT systems with respect to their fine-tuning method (FT, unifPE or SHAPE), backbone model (e.g. NLLB) and training corpus (U for TED-U or G for TED-G. More precisely, we denote MT systems trained on TED-U (resp. TED-G) as FT-NLLB-U (resp. FT-NLLB-G), Unif-NLLB-U (resp. Unif-NLLB-G) when fine-tuning with unifPE, and SHAPE-NLLB-U (resp. SHAPE-NLLB-G).
We also experiment with an LLM-based architecture, TowerBase-7B [^16] https://huggingface.co/Unbabel/TowerBase-7B-v0.1 (Alves et al., 2024) (TowerBase for short), derived from Llama2 (Touvron et al., 2023) using translation-related tasks. TowerBase uses RoPE (Su et al., 2024) to encode RPEs. As mentioned by Peng et al. (2024), they nonetheless encode some form of APE signal in some dimensions, and may therefore be also mildly impacted by the PE training distribution. We refer to the models based on TowerBase as FT-Tower-U (resp. FT-Tower-G), Unif-Tower-U (resp. Unif-Tower-G) and SHAPE-Tower-U (resp. SHAPE-Tower-G) the model fine-tuned on TED-U (resp. TED-G) with original PEs, unifPE or SHAPE.
Both backbone models were pretrained with large amounts of EN–FR data; we focus exclusively on the EN into FR direction. disabledisabletodo: disableCheck directionDetails on fine-tuning and decoding parameters can be found in Appendix A.4.
@@ 6.1 Length Issues
We report the ds-BLEU (Table 2) and COMET (Appendix, Table 7) scores of the pretrained models NLLB and TowerBase for multiple test sets, varying the average input segment lengths from one sentence to the maximum input length used in training. [^17] As explained in Section 3.2, these COMET scores require the realignment of target sentences with the reference.disabledisabletodo: disableCOMET scores in appendix ? For NLLB, we observe a drop of around $10$ ds-BLEU points and about $0.2$ COMET points when translating test sets of $l_{max}=256$ instead of isolated sentences. Scores and their associated brevity penalties (BPs) only get worse with larger context lengths. For TowerBase, the decrease in BLEU is more progressive, with a sharp decline for all test sets for $l_{max}>1024$. The related COMET scores plummet immediately with a context size of $256$. Even though TowerBase is based on Llama2, which accepts inputs up to $4096$ tokens, the continued pretraining that was used mostly uses isolated sentences, which introduces an inductive bias affecting its ability to translate long texts.
@@
TED-U TED-G FT Unif SHAPE
 FT vs Unif  FT vs SHAPE  FT vs Unif FT vs SHAPE U vs G U vs G U vs G
sent 3.3 (0.00) 4.0 (0.00) 1.2 (0.00) 2.4 (0.00) - -2.1 (0.00) -1.6 (0.00)
256 - 0.7 (0.00) - - - -0.6 (0.01) -0.7 (0.00)
512 -0.5 (0.01) 1.7 (0.01) - 2.3 (0.00) -0.7 (0.00) -0.4 (0.01) -
768 -0.8 (0.00) 2.7 (0.00) -3.7 (0.00) - 5.5 (0.00) 2.6 (0.00) 4.5 (0.00)
1024 - 1.6 (0.00) -7.8 (0.00) -1.8 (0.04) 12.2 (0.00) 5.1 (0.00) 8.8 (0.00)
1200 -2.3 (0.00) 2.3 (0.02) -8.5 (0.00) - 12.7 (0.00) 6.5 (0.00) 8.8 (0.00)
1600 -2.6 (0.01) - -8.8 (0.00) -2.6 (0.01) 11.8 (0.00) 5.6 (0.00) 8.3 (0.00)
2048 -3.3 (0.00) - -7.3 (0.00) - 10.7 (0.00) 6.8 (0.00) 11.3 (0.00)
sent 1.9 (0.00) 2.9 (0.00) 0.9 (0.00) 1.4 (0.00) - -0.9 (0.00) -1.4 (0.00)
256 - 0.8 (0.00) 0.4 (0.00) 0.7 (0.00) -0.8 (0.00) -0.5 (0.01) -0.9 (0.00)
512 -0.6 (0.02) 2.9 (0.00) - 4.3 (0.00) -0.4 (0.03) - -
768 -0.7 (0.00) 4.7 (0.00) -4.7 (0.00) - 11.2 (0.00) 7.2 (0.00) 7.3 (0.00)
1024 -2.2 (0.00) 3.3 (0.00) -7.5 (0.00) -1.8 (0.02) 19.3 (0.00) 14.0 (0.00) 14.2 (0.00)
1200 -5.3 (0.00) 2.7 (0.02) -6.5 (0.00) - 15.7 (0.00) 14.5 (0.00) 11.9 (0.00)
1600 - - -6.4 (0.00) - 11.5 (0.00) 6.6 (0.00) 9.2 (0.00)
2048 - 2.1 (0.04) -4.7 (0.00) - 10.4 (0.00) 7.2 (0.00) 8.3 (0.00)
Table 4: Average difference (and p-values) in ds-BLEU (top) evaluated on full TED talks and 100×100$\times$100 ×COMET (bottom) evaluated on realigned sentences for NLLB. Left and middle: paired comparison between fine-tuning with the original PEs (FT), unifPE (Unif) and SHAPE on TED-U and TED-G respectively. Right: differences between fine-tuning on TED-U (U) and TED-G (G). - for p-values $>0.05$. Bold values when the two metrics disagree on significativity.
@@
Again using the paired comparison methodology, we compare the performance of DLFT with original PEs, unifPE and SHAPE. As shown in the left and middle parts of Table 4, fine-tuning using unifPE leads to steady improvements in translation scores for all test lengths, especially for systems fine-tuned with the unbalanced corpus (TED-G). The only exception is for sentence-level translations, which remain marginally better using standard DLFT than with unifPE. In contrast, SHAPE only improves DLFT performance on the TED-G corpus and for translation windows greater than $1024$ tokens, due to the under-representation of small position indices during training, as shown in Figure 1. As Tables 11, 12, 13, 14 and 15 show, these improvements remain moderate, and the length issues continue to strongly impact translation scores, especially for test documents of $1024$ tokens or more.
NLLB FT-U Unif-U SHAPE-U FT-G Unif-G SHAPE-G
$p_{0}$-$p_{1}$ 12.7 - - 4.0 1.6 3.0 5.0
$p_{1}$-$p_{2}$ 7.2 - - -2.0 1.9 2.4 -
$p_{2}$-$p_{3}$ 2.9 1.0 -1.3 - 2.4 - -2.1
$p_{3}$-$p_{4}$ 7.2 5.5 4.6 7.3 26.1 10.7 13.9
$p_{4}$-$p_{5}$ 3.9 - - 8.3 4.7 4.3
$p_{5}$-$p_{6}$ 3.3 31.6 27.1 19.5 6.1 15.4 15.3
Tower FT-U Unif-U SHAPE-U FT-G Unif-G SHAPE-G
$p_{0}$-$p_{1}$ 1.2 - - - - - -
$p_{1}$-$p_{2}$ 0.6 0.6 -
$p_{2}$-$p_{3}$
$p_{3}$-$p_{4}$ 4.9 1.1 0.6 1.4 1.1 1.3 1.1
$p_{4}$-$p_{5}$ 1.7 1.4 2.1 1.7 2.0 1.8 2.2
$p_{5}$-$p_{6}$ 26.3 24.5 26.0 25.2 27.1 26.4 27.0
Table 5: Average difference of 100×100$\times$100 ×COMET-score evaluated on $794$ sentence pairs, translated at different positions (e.g. $p_{0}$ and $p_{1}$ with $p_{0}<p_{1}$) by NLLB-based systems (top) and TowerBase-based systems (bottom). Olive text for p-values $>0.01$. - for p-values $>0.05$.
For TowerBase, unifPE does not yield any significant difference with standard DLFT, and SHAPE occasionally delivers slight improvements (see Appendix, Table 16), likely because this model relies on RPEs. From these comparisons, we conclude that unifPE partly resolves length issues for nllb, but hardly changes the situation for TowerBase.
@@
In the right part of Table 4, we evaluate the impact of the length distribution during fine-tuning for NLLB: the balanced distribution (TED-U) slightly but consistently underperforms the use of TED-G for short documents (fewer than $512$ tokens), a trend that is reversed for longer documents with strong improvement (over $768$ tokens). Manipulating the distribution of PEs with unifPE reduces the gap between the two fine-tuning corpora and makes the model more robust to document lengths rarely observed (or even unobserved) during fine-tuning. This analysis again reveals small differences between using ds-BLEU and COMET scores: in nine cases out of $56$ comparisons (marked in bold), one metric detects a difference that is non-significant for the other.
@@ 6.2 Position Bias
To investigate potential translation issues related to large position indices, we collected the $794$ sentences that come from the final part of long talks and for which varying the window length also varied the position index. For each of them, we have seven translations, corresponding to positions $\{p_{0}^{j},\dots,p_{6}^{j}\},j\in\{1,\dots,794\}$. The average values for $\{p_{0}^{j},\dots,p_{6}^{j}\}$ are $\{p_{0},\dots,p_{6}\}=[66,173,262,335,585,779,1477]$. For this subset of sentences, we performed a paired t-test to compare the impact of the position on the translation score (using COMET as the only metric). We observe in Table 5 that in almost all comparisons but three, a small position index is preferable to a larger one. This suggests that one of the main challenges faced by Doc2Doc with large context lengths is to control the quality degradation for the final parts of the input text. Here again, unifPE slightly mitigates the problem for NLLB models compared with original PEs and SHAPE, but no such improvement is observed for TowerBase.disabledisabletodo: disableWhy is tower different?
256 512 768 1024 1200 1600 2048
NLLB 0.04 0.35 0.49 0.66 0.64 0.74 0.81
FT-U 0.01 0.03 0.08 0.09 0.13 0.26 0.44
Unif-U 0.01 0.03 0.07 0.10 0.20 0.34 0.32
SHAPE-U 0.03 0.08 0.11 0.20 0.36 0.39 0.46
FT-G 0.01 0.03 0.11 0.31 0.40 0.57 0.69
Unif-G 0.01 0.04 0.16 0.20 0.27 0.25 0.36
SHAPE-G 0.02 0.05 0.12 0.17 0.21 0.24 0.28
256 512 768 1024 1200 1600 2048
Tower 0.01 0.05 0.13 0.30 0.29 0.45 0.64
FT-U 0.01 0.05 0.10 0.15 0.13 0.23 0.59
Unif-U 0.02 0.05 0.09 0.15 0.16 0.28 0.61
SHAPE-U 0.02 0.05 0.08 0.15 0.17 0.21 0.59
FT-G 0.02 0.05 0.10 0.15 0.19 0.26 0.62
Unif-G 0.02 0.07 0.10 0.16 0.20 0.28 0.64
SHAPE-G 0.02 0.05 0.07 0.17 0.17 0.27 0.62
@@ 6.3 Repeated $n$-grams in Translation
One obvious problem with holistic translations produced by Nllb is the generation of outputs that are too short. A closer look at translation outputs also reveals that outputs contain many instances of repeated texts, usually occurring in the final part of the translation. To quantify this problem, we compute the percentage of translations of pseudo-documents in which the repetition of a long $n$-gram (with $n\geq 10$) is detected. Detailed results are given in Table 6. For all systems and fine-tuning strategies, the percentage of repetitions increases with the length, a problem that seems (for large text lengths) slightly more severe for TowerBase, which has a much better BP, than Nllb.
@@ 7 Conclusion
In this work, we have studied the ability of current MT architectures to handle long input texts, ideally entire documents, and to translate them holistically. Our analyses are based on systematic comparisons of translation outputs computed with varying input lengths, which are then evaluated with two automatic metrics. They consistently show that, even when the test document lengths match that of the training set and remain within the model limits, the translation scores tend to decrease with the source length, a degradation that mostly impacts sentences occurring far from the beginning of the document. We also show that manipulating the training distribution of lengths or PEs has a positive effect for APE-based models, which vanishes in RoPE-based models like TowerBase.disabledisabletodo: disableMake sure RPE is defined These results finally confirm the robustness of sentence-level baselines. They hint at the need to improve existing models to truly benefit from the potential of document-level MT, for instance by constraining the attention mechanism to simulate a form of sentence alignment, by improving the memorization capacities of existing architectures, or by ensuring that the generation algorithm does not eventually get trapped in repetition loops. These are some of the directions we wish to explore in future work.disabledisabletodo: disablePerhaps: more citations for this disable,color=cyan!30disable,color=cyan!30todo: disable,color=cyan!30Future: the errors beyond repetitions and missing material disabledisabletodo: disablethere should not be TEDtalks in NLLB training set, as the NLLB papers used IWSLT TED talks as benchmark to show the model performance, where is the only mention os TED talks.
@@ 8 Limitations
The empirical observations reported in this paper are based on one single language direction,disabledisabletodo: disablecheck that this is clear and one domain (TEDtalks). This experimental design is motivated by (a) the fact that French-English is considered an easy pair for MT, with large sets of parallel training data; (b) TEDtalks data are a standard benchmark for document-level MT, and crucially contain very long parallel documents, allowing us to implement our evaluation methodology on a large range of length values. Furthermore, these datasets are not included in the training data of our models. We contend that the length issues observed in these favorable conditions for two representative systems would only be worse for more difficult or less-resourced language pairs.
@@ 10 Carbon Impact Statement
The experiments were conducted on a private infrastructure using a single A100 SXM4 GPU, with a carbon efficiency of $0.432$ kgCO2eq/kWh. The average time required for fine-tuning and checkpoint selection was $14.21$ hours for the six NLLB models, and $5.6$ hours for the Towerbase models. The average emissions are estimated to be $2.45$ kgCO2eq for NLLB-based models and $0.97$ kgCO2eq for models derived from TowerBase, with no offset applied. These estimations were based on the Machine Learning Impact calculator [^18] https://mlco2.github.io/impact#compute (Lacoste et al., 2019).
Acknowledgments
This work was supported by the French national agency ANR as part of the MaTOS project. [^19] http://anr-matos.fr/ Rachel Bawden was also partly funded by her chair position in the PRAIRIE institute funded by ANR as part of the “Investissements d’avenir” programme under reference ANR19-P3IA-0001. The authors are grateful to the anonymous reviewers for their insightful comments and suggestions and to Paul Lerner for his review and feedback on a preliminary draft of this work.